OpenStudy (anonymous):

find the zeros of y= x^3+6x^2-7x

OpenStudy (anonymous):

\[ x^3+6x^2-7x=0\] \[x(x^2+6x-7)=0\] is a start

OpenStudy (anonymous):

then factor further as \[x(x+7)(x-1)=0\] then it should be ok